OpenStudy (anonymous):

f(x)=2x^2+7x+3 Use grouping method Use quadratic method Use completing the square method

OpenStudy (whpalmer4):

Do you know how to do any of these?

OpenStudy (anonymous):

No this is a pre-test @whpalmer4

OpenStudy (anonymous):

But can you help me?

OpenStudy (anonymous):

@Hero @whpalmer4 ?

hero (hero):

As an example, Suppose f(x) = x^2 + 5x + 6

hero (hero):

Then, to factor by grouping, find two numbers m and n so that m × n = 6 m + n = 5 Notice that 2 × 3 = 6 and 2 + 3 = 5 Basically, now, since we have found the values of m and n, then substitute 5 with 2 + 3: x^2 + 5x + 6 = x^2 + (2 + 3)x + 6 = x^2 + 2x + 3x + 6 Now factor the first two terms, then factor the last two terms. = x(x + 2) + 3(x + 2) Now factor the remaining term (x + 2) that is common to both factorizations: = (x + 2)(x + 3) So x^2 + 5x + 6 = (x + 2)(x + 3)

hero (hero):

To factor by quadratic equation \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a}\) Let a = 1, b = 5 and c = 6 Then \(x = \dfrac{-5 \pm \sqrt{5^2 - 4(1)(6)}}{2(1)}\) \(x = \dfrac{-5 \pm \sqrt{25 - 24}}{2}\) \(x = \dfrac{-5 \pm \sqrt{1}}{2}\) \(x = \dfrac{-5 \pm 1}{2}\) \(x = \dfrac{\pm 1 - 5}{2}\) \(x = \pm \dfrac{1 - 5}{2}\) \(x = - \dfrac{1 + 5}{2} \text{or} \dfrac{1 - 5}{2}\) \(x = -\dfrac{6}{2} \text{or} \dfrac{-4}{2}\) \(x = -3 \space \text{or} -2\)

hero (hero):

Then x = -3 x + 3 = 0 or x = -2 x + 2 = 0 and (x + 3)(x + 2) = 0 by zero product property

hero (hero):

For complete the square: Let f(x) = 0 Then x^2 + 5x + 6 = 0 Subtract 6 from both sides: x^2 + 5x = -6 Then add \(\left(\dfrac{b}{2}\right)^2\) to both sides Since b = 5, then \(\left(\dfrac{b}{2}\right)^2 = \dfrac{5^2}{4} = \dfrac{25}{4}\) So add \(\dfrac{25}{4}\) to both sides: \(x^2 + 5x + \dfrac{25}{4} = \dfrac{25}{4} - 6\) \(x^2 + 5x + \dfrac{25}{4} = \dfrac{25}{4} - \dfrac{24}{4}\) \(x^2 + 5x + \dfrac{25}{4} = \dfrac{1}{4}\) Remember that in order to express the right side as a binomial square we need mn = 25/4 m + n = 5 In this case m = n = 5/2 Since (5/2)(5/2) = 25/4 and 5/2 + 5/2 = 10/2 = 5 So \(x^2 + 5x + \dfrac{25}{4} = \left(x + \dfrac{5}{2}\right)^2\) and \(\left(x + \dfrac{5}{2}\right)^2 = \dfrac{1}{4}\) Take the square root of both sides: \(x + \dfrac{5}{2} = \pm\dfrac{1}{2}\) Isolate x: \(x = \pm\dfrac{1}{2} - \dfrac{5}{2}\) \(x = -2,\) \(x = -3\) \(x + 2 = 0\) \(x + 3 = 0\) \((x + 2)(x + 3) = 0\) \((x + 2)(x + 3) = f(x)\)

OpenStudy (anonymous):

Man thank you so much @hero helps so much omg

hero (hero):

yw